OpenStudy (anonymous):

will give medal a number raised to a negative exponent is negative. (1 point) a. always b. never c. sometimes

OpenStudy (anonymous):

sometimes.

OpenStudy (anonymous):

sometimes

OpenStudy (anonymous):

sometimes

OpenStudy (anonymous):

sometimes

OpenStudy (anonymous):

thanks everyone, comment on any question of mine and ill give you all medals :)

OpenStudy (anonymous):

let me explaine. Say, our number can be positive or negative or zero. Lets forget zero, because zero to some power doesnot make sense. case 1. illustration of positive; say 1 raised to power -1. ie; 1^(-1) =1/(1^1)=1. positive. But then, if our number is (-1), and raised to -1, ie; (-1)^(-1)= 1/((-1)^1=-1.

OpenStudy (gorv):

If a negative number is raised to an odd negative exponent, the answer will still be negative. --> Sometimes